OpenStudy (anonymous):

find the point of intersection of y=e^2x and y =e^X+2

OpenStudy (anonymous):

you should have taken one equation as z or anything else both can't be y

OpenStudy (anonymous):

huh? thats what the question says

OpenStudy (anonymous):

so i was thinking you can say that e^2x = e^X+2 but i don't now how to solve from there onwards

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is capital x given? or is that normal x?

OpenStudy (anonymous):

oh sorry its meant to be a normal x LOL

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

e^2x-e^x=2 e^x(e^x - 1) = 2 so either e^x = 2 or e^x = 3

OpenStudy (anonymous):

so x = log(2) ~ .7

OpenStudy (anonymous):

or x = log(3) ~ .47

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

no no my dear this is not right.

OpenStudy (anonymous):

well i dont know how is the log(3) coming here

OpenStudy (anonymous):

thank you mridul1 anyways!

OpenStudy (anonymous):

@vertus i answered in previous post @mridul1 if you have \[x(x-1)=2\] you cannot conclude that \[x=2\] or \[x=3\] or anything for that matter. you have to solve the quadratic \[x^2-x-2=0\]

OpenStudy (anonymous):

oh yes you're right my silly mistake thanks for pointing it out

OpenStudy (anonymous):

^.^